OpenStudy (zepp):

Derivative of a constant is 0!? o.O How come?

OpenStudy (agent47):

A constant function is just a horizontal line. Slope of a horizontal line is 0.

OpenStudy (agent47):

And since the slope of a function at a point is a derivative at that point, derivative of any constant is 0

OpenStudy (zepp):

Oh I see, thanks :)

OpenStudy (anonymous):

Derivative represents the small change in something.. As constant is a value that remains same, so there is 0 change in the constant...

OpenStudy (saifoo.khan):

Agent47 earned a fan.

OpenStudy (zepp):

Got it, thanks guys! :)

OpenStudy (anonymous):

if you want to be silly you can prove it directly from the difference quotient let \(f(x)=c\) then \[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}=\lim_{x\to 0}\frac{c-c}{0}=0\]

OpenStudy (saifoo.khan):

@zepp , your fav method^

OpenStudy (anonymous):

typo \[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}=\lim_{h\to 0}\frac{c-c}{h}=0\]

OpenStudy (zepp):

You know me so well, saif ;D Thanks @satellite73!